OpenStudy (anonymous):

Solve: 7 – 5(2 – x) = –(x – 8)

OpenStudy (anonymous):

so to start off can someone explain to me how I would start to solve this equation

OpenStudy (anonymous):

Distribute -5 into the parenthesis. -5 * 2 = ? -5 * -x = ?

OpenStudy (swamijatin2):

(7-5) x (2-5)=-(x-8)

OpenStudy (swamijatin2):

now , -6=-x+8 -6-8=-x -14=-x x=14

OpenStudy (anonymous):

ok so that would be 3*-3=(-8) which is -6-8=-14 thank you swamijatin

OpenStudy (swamijatin2):

yes

OpenStudy (anonymous):

A. B. C. –4 D. A. 5/4 B.7/2 C. –4 D. 11/6

OpenStudy (anonymous):

so that is the answers it gives and im totally confused

OpenStudy (swamijatin2):

ples wait

OpenStudy (anonymous):

ok

OpenStudy (swamijatin2):

7-5(2-x)=-(x-8) 7-10+5x=-x+8 6x=11 x=11/6

OpenStudy (swamijatin2):

d.

OpenStudy (anonymous):

oh ok I see what you did that makes since thank you again